OpenStudy (anonymous):

what ordered paris are a solution of x+1/2y=1?

OpenStudy (solomonzelman):

\[x+\frac{1}{2}y=1~~~~~~~~~~~~~~~~~~~this?\]

OpenStudy (anonymous):

yess

OpenStudy (solomonzelman):

\[2x+y=2\]

OpenStudy (solomonzelman):

so x=1 2(1)+y=2 2+y=2 y=0 so first point is (1,0) x=2 2(2)+y=2 4+y=2 y=-2 so first point is (2,-2) see?